Truut, Ina prepared the following production repcrt FlUn the missrg amourts. (Round your "Cost per Equlvalent Unit" answers to 2 decimal places.) TROUT, I Process Costing Production Raport (Walghted Average Mothad) INITS Physical 9 Accounted for Units Work in 3,000 nto Procuction Tutal Units 7.00 Sten 2 Equivalent Units Units Accounted far as Materials Comversion dand Transferre Wark in Process ntal Lnito 000 7,000 5,500 Step 3: Cost per Equivalent Uni Materials Comversion COST Total Cost Cost of Beginring Inventory $18,000 S 14.000 32000 122,000 str Total Cost of Urits in S 220 000 5 500 12.500 Probet Uts Equvalent Uni Cost 40.00 Step 4: Reconcile Work in Materials Comverrio Total 14.000 122.000 206 0001 328 000 Cost to be Accaunted far Cost Begning Wark in PrObes ind Cords S18,000 Addtonal information: Trout adds direct materials at the beginning af the process, while panversian costs are incurred untarmly thraughout t prucess. At the beginning of the penod, Work in Process Inventory was 40 % pamplete. At the end af the period, Work in Process Inventory was 70%% Damplete.
Truut, Inc. prepared the folowing praduction repart Required: Fil in the missing amounts. (Round your "Cost per Equivalent Unit" answers to 2 decimal places.) Equivalent Units Accounted for as Fellkrax Materials Conversion and Tranatorrad 2000 5,000 5,000 7000 7000 2000 Ending V l Units k in Process 5,500 Equivalent Step 3: Cost per COSTS Materials Conversion Total , Cost of Baginning mat Prriod Casts $18,000 14.000 000 122.000 206.000 328.000 r 220,000 00 7,00 Fn oUnite in 12.500 Eu nniert Unit Work in Step 4: Reco Procaze Cost to ba Accountad for Materials Conersion Total Po WWark in Current Period Costs $18,000 14,000 0on 122,000 206,000 328,000 Total Costs 140.000220,000 000 Cast Accourted for as Comtieled andi S Transtorad Ending Inwentory 140.ar 220.000 360,000 ... Additional information: Trout adds direct materials at the beginning of the process, whie conversion costs are incurred uniformly thraughout the process. At the beginning of the period, Work in Process Imventory was 40 % complete. At the end of the period, Wark in Process Inventary was 70 % complete.
Answer #1

CORRECTED FILL IN THE BLANKS :

Process costing production report (WAM):
Units to be Accounted for Physical Units
Beginning WIP 3000
Started into Produ. 4000
Total Units 7000
Units Accounted for as follows: Physical Units Materials Conversion costs
Beginning WIP (under WAM) 3000 2000 2000 (being under WAM)
Started and completed 0 0 0
Completed and Transfere 2000
Ending WIP 5000 5000 3500 (70%Complete)
Equivalent Units 7000 7000 5500
Cost per Equivalent unit:
Cost of Beg WIP 18000 14000
Current period Costs 122000 206000
total cost of units in process 140000 220000
Equivalent units 7000 5500
cost per Eq. unit 20 40
Reconcile WIP :
Cost to be Accounted for: Materials Conversion Total Costs
Beg WIP 18000 14000 32000
Current period costs 122000 206000 328000
Total Costs 140000 220000 360000
Cost Accounted for:
Units completed and Transferred 2000*20=40000 2000*40=80000 120000
Ending Inventory 5000*20=100000 3500*40=140000 240000
Total Costs accounted for 140000 220000 360000